5x – 3y = -6
3. Describe the steps necessary to graph this equation with the x- and y-intercepts. You have to
4. Describe how to graph this equation by using the slope and y-intercept.
2. Describe how to graph this equation by plotting points from a table of values.
1. Is this a linear equation? Why? What are the characteristics of this equation that help you to identify its type

Answer by stanbon(75887)

5x – 3y = -6
3. Describe the steps necessary to graph this equation with the x- and y-intercepts.
If x = 0, -3y=-6, y = 2
If y = 0, 5x =-6, x = -6/5
Plot (0,2) and (-6/5,0) and draw a line thru them
-------------------------
4. Describe how to graph this equation by using the slope and y-intercept.
y = (5x+6)/3
y = (5/3)x + 2
slope = 5/3 ; y-int = (0,2)
Plot (0,2)
slope mean y increases by 5 when x increases by 3
So, plot (0+3,2+5) = (3,7)
Draw a line thru (0,2),(3,7)
----------------------------------------
2. Describe how to graph this equation by plotting points from a table of values.
y = (5/3)x + 2
If x = 3, y = 5 + 2 = 7
If x = 6, y = 5*2+2 = 12
Plot (3,7) and (6,12) and draw a line thru them.
----------------
1. Is this a linear equation? Yes
Why? What are the characteristics of this equation that help you to identify its type
Ans:: slope-intercept form
-------------
Cheers,
Stan H.
----------